# OPeRA: A Dataset of Observation, Persona, Rationale, and Action for Evaluating LLMs on Human Online Shopping Behavior Simulation

### v2

@@ Appendix F Segmentation Interval Threshold Justification
Appendix G Dataset Statistics
Table  6 and  3 show the statistics for processed dataset.
Processed Dataset
@@ Appendix G Dataset Statistics
Purchase 320 (3.28%) 22 (7.91%) 9.50
Text Input 660 (6.77%) 0 –
Termination 395 (4.05%) 0 –
@@
Table 6: Processed dataset statistics. “Len” reports the average rationale length (in words) for each action type.
Appendix H Experiment Prompt Design
@@
Click on a button or clickable element identified by name or selector‘, its further classified with click_type including search‘, filter‘, other‘.
@@
"click_type": "click_type",
"selector": "selector"
@@
#### 4. purchase‘:
Indicate intent to purchase by clicking on any purchase-related button. This includes clicking "Add to Cart", "Buy Now", "Set Subscription", or "Checkout" buttons. Any of these actions are considered as showing purchase intention.
"type": "purchase"
@@
PROPMT_FOR_RATIONALE_ACTION_PREDICTION="""
<IMPORTANT>
You are simulating a real shopper using Amazon.com.
Your task is to generate a *short rationale (several simple words)* that explains why the user takes a next shopping action.
You need to pretend that you are a real user shopping on amazon.com.
The history action, rationale and context will be provided to you.
The rationale should reflect a natural, human-like thought process behind the action.
If there are previous rationales, you should follow the style of the previous rationales.
Use short, realistic sentences. Example:
- "I saw a discount."
- "I want to compare more options."
- "Im looking for something with better reviews."
</IMPORTANT>
Same as above ...
# Output Format
You need to predict the next action AND rationale. Your output should follow a strict JSON format:
"rationale": "<rationale>" // rationale goes here, a string
"action": {
// action goes here
"type": "<type>",
<IMPORTANT>
OUTPUT A SINGLE JSON OBJECT, NOTHING ELSE.
</IMPORTANT>

### v4

@@ Appendix F Segmentation Interval Threshold Justification
Appendix G Click Type Description
This section provides detailed descriptions for the click type categories used in the OPeRA dataset, as presented in Table 3.2. The following categories were established through empirical analysis of user click action distributions.
review (20.8%): Clicks on review-related elements, including review images, star ratings, review filters, etc.
search (15.1%): Clicks on search button or search box.
product_option (13.9%): Clicks on product options such as size selectors, color options etc.
product_link (10.6%): Clicks on product images, product titles, or product links that navigate users to product detail pages.
other (8.9%): Miscellaneous clicks that do not fall into the above predefined categories.
purchase (6.4%): Clicks on purchase-intention elements including “Add to Cart”, “Buy Now”, “Subscribe”, and “Checkout”.
nav_bar (5.6%): Clicks on navigation bar elements such as category menus, amazon logo etc.
page_related (3.9%): Clicks on pagination controls, carousel navigation buttons that control page content display.
quantity (3.8%): Clicks on quantity adjustment controls including increase/decrease buttons and item deletion buttons.
suggested_term (3.6%): Clicks on search suggestions.
cart_side_bar (2.9%): Clicks on elements in shopping cart sidebar.
cart_page_select (2.8%): Clicks on selection elements such as item checkbox in the cart page.
filter (1.8%): Clicks on filtering elements including price filters, brand filters, rating filters, and other product refinement controls.
Appendix H Next Action Prediction Experiment Results
Table 6 shows the full results of next action prediction task.
@@ Appendix H Next Action Prediction Experiment Results
  w/o persona 22.06 82.32 45.55 43.45 55.55 58.47
   w/o rationale 21.28 83.13 34.93 42.63 53.33 51.17
@@
  w/o persona 10.75 82.28 25.33 22.76 50.00 43.10
   w/o rationale 10.08 81.08 26.06 20.29 47.78 43.10
[…]
  w/o persona 8.31 78.59 23.69 18.59 28.89 33.21
   w/o rationale 8.76 80.23 23.60 19.22 31.11 34.19
[…]
Table 6: Evaluation of actions in next action prediction task. “Claude-3.7”: Claude-3.7-Sonnet, “Llama-3.3”: Llama-3.3-70B-Instruct. All metrics are reported as percentages (%). Instance size $n=902$.
Appendix I Experiment Prompt Design
@@
Click on a button or clickable element identified by name‘.